# Burer-Monteiro factorizability of nuclear norm regularized optimization

### v1

@@ Appendix A Pseudo-stationarity
    where we used the fact that $x_{i}$ and $y_{\tau(i)}$ are 00. Notice that the objective of the above problem is $-(g_{i}-v_{\tau(i)})^{2}$. Clearly, the optimal value of (B.30) is 00, and is achieved if and only if
    $g_{i}=v_{\tau(i)}\in[0,1].$ (B.31)
Consequently, by the solution sets given in (B.22), (B.25), (B.28), (B.29) and (B.31), we know the solution set of (B.16) is nonempty. The optimal value is obtained by summing all the optimal values given in the four cases. Moreover, every solution $(\bar{x},\bar{g},\bar{y},\bar{v})$ of (B.16) satisfies $\|\bar{g}\|_{\infty}>1$ if and only if $|\mathcal{J}_{3}^{\tau}|>0$, according to the structure of $\bar{g}$ given in (B.28), (B.29) and (B.31). ∎
Proposition B.4.
Let $r^{*}\in[m]\cup\{0\},r\in[m]$, and $\infty>L\geq\mu>0$. Let $G$ be the objective function of B.15 and let $\kappa:=\frac{L}{\mu}\geq 1$. If $r^{*}$, $r$ and $\kappa$ satisfy any of the following conditions, then there is no feasible $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ to (B.15) satisfying $\|\bar{g}\|_{\infty}>\lambda$ and $G(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})\geq 0$.
  1. 1.
    $r^{*}=0$ or $r=m$.
  2. 2.
    $r>r^{*}$ and $\kappa=3$.
  3. 3.
    $r\geq r^{*}$ and $\kappa<3$.
  4. 4.
    $r\geq r^{*}$, $\kappa>3$ and $\frac{r-r^{*}}{1-\frac{4}{(\kappa-1)^{2}}}>\min\{r,m-r^{*}\}$.
@@
By the change of variables $(x,g,y,v,\tau,w)\leftarrow(x/\lambda,g/\lambda,y/\lambda,v/\lambda,\tau,w/\lambda)$, we see that B.15 can be reduced to the following problem:
$\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m},w\in\mathbb{R}\end{subarray}}$ $\lambda^{2}\bigg{[}\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (B.32)
@@
Since dropping the constant $\lambda^{2}$ won’t affect the sign of the function value and our claim only concerns the feasible set of (B.15) and the sign of its objective value, we shall consider the following optimization problem instead:
$\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m},w\in\mathbb{R}\end{subarray}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (B.33)
@@
Notice that when $\tau$ and $w$ are fixed, B.33 becomes B.16. Applying Lemma B.3, setting $d_{\tau}=|\mathcal{J}^{\tau}_{2}|$ and recalling the definition of $\mathcal{J}^{\tau}_{3}$ in (B.14), we see that the solution set $\Omega_{w,\tau}$ of (B.16) is nonempty, and for all $(\bar{x},\bar{g},\bar{y},\bar{v})\in\Omega_{w,\tau}$, it holds that $\|\bar{g}\|_{\infty}>1$ if and only if $r^{*}-r+d_{\tau}>0$.
@@
$H(d,w):=\left(-L\mu d+(r^{*}-r+d)\frac{L(L-\mu)^{2}}{4\mu}\right)w^{2}.$ (B.34)
Then in view of (B.14) and (B.17), the optimal value of (B.16) is given by $H(|\mathcal{J}^{\tau}_{2}|,w)$. Moreover, we see that (B.33) is equivalent to the following problem
$\sup_{w\in\mathbb{R},~{}d\in\mathbb{N}_{0}}H(d,w)\quad s.t.~{}w>0,~{}r-r^{*}\leq d\leq\min\{r,m-r^{*}\},$ (B.35)
where the bound for $d$ comes from the requirement that $|\mathcal{J}_{i}^{\tau}|\geq 0$ for $i\in[4]$ in B.14.
@@
    The optimal value of (B.35) is nonpositive, and (B.35) has no feasible solution $(d,w)$ satisfying $H(d,w)\geq 0$ and $r^{*}-r+d>0$.
    In this scenario, we claim that (B.15) has no feasible solution $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ with $\|\bar{g}\|_{\infty}>\lambda$ and $G(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})\geq 0$, where $G$ is the objective of B.15. A short proof is provided below.
    Suppose such a feasible solution $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ of (B.15) exists, then either $(\bar{x},\bar{g},\bar{y},\bar{v})/\lambda$ is optimal for (B.16) with $w=\bar{w}/\lambda$ and $\tau=\bar{\tau}$, or $(\bar{x},\bar{g},\bar{y},\bar{v})/\lambda$ is not optimal. In the latter case, the optimal value of (B.35) must be positive. In the former case, we see from Lemma B.3 that $|\mathcal{J}_{3}^{\bar{\tau}}|>0$, and hence (B.35) has a feasible solution $(\tilde{d},\tilde{w})=(|\mathcal{J}^{\bar{\tau}}_{2}|,\bar{w}/\lambda)$ with $H(\tilde{d},\tilde{w})\geq 0$ and $r^{*}-r+\tilde{d}=r^{*}-r+|\mathcal{J}^{\bar{\tau}}_{2}|=|\mathcal{J}^{\bar{\tau}}_{3}|>0$ (see (B.14)). Both cases yield a contradiction.
@@
    There exists a feasible solution $(d,\tilde{w})$ of B.35 satisfying $H(d,\tilde{w})\geq 0$ and $r^{*}-r+d>0$.
    In this scenario, (B.15) has a feasible solution $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ with $G(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})\geq 0$ and $\|\bar{g}\|_{\infty}>\lambda$. Indeed, we just need to take $\bar{\tau}\in\mathfrak{P}_{m}$ satisfying $|\mathcal{J}_{2}^{\bar{\tau}}|=d$, and then take $(\tilde{x},\tilde{g},\tilde{y},\tilde{v})$ to be the optimal solution of (B.16) with $\tau=\bar{\tau}$ and $w=\tilde{w}$, and set $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{w})=\lambda(\tilde{x},\tilde{g},\tilde{y},\tilde{v},\tilde{w})$.
We note that the classification in (S1) and (S2) is not complete, since we cannot say anything if the optimal value of (B.35) is positive and there is no feasible solution $(d,w)$ of (B.35) satisfying $H(d,w)\geq 0$ and $r^{*}-r+d>0$. Nevertheless, the two scenarios in (S1) and (S2) are enough for our proof. Consider the following cases on $r$, $r^{*}$ and $\kappa:=L/\mu$.
@@
    $r=m$ or $r^{*}=0$. If $r=m$, then every feasible point $(x,g,y,v,\tau,w)$ of (B.12) satisfies $\|g\|_{\infty}=\lambda$; moreover, every feasible point $(d,w)$ to (B.35) must satisfy $d=r-r^{*}$ and $H(d,w)=-L\mu dw^{2}\leq 0$. Then (S1) holds. If $r^{*}=0$, we see that every feasible point $(d,w)$ to (B.35) must satisfy $d=r$. Then, in view of (B.34), we can rewrite (B.35) as:
@@
    This means that every feasible solution of (B.35) has a negative objective function value. Then (S1) holds.
@@
    Then the optimal value of (B.35) is 00, and for all feasible solution $(d,w)$ of (B.35) with $r^{*}-r+d>0$ it holds that $d>r-r^{*}\geq 0$, and hence $H(d,w)<0$. Then (S1) holds.
@@
    $m>r\geq r^{*}>0$ and $L>\mu$ (i.e., $\kappa>1$). If $\kappa=\frac{L}{\mu}=3$, then we have $-L\mu+\frac{L(L-\mu)^{2}}{4\mu}=L\mu(\frac{(\kappa-1)^{2}}{4}-1)=0$. Therefore, (B.35) is equivalent to that:
@@
    In this case, we clearly see that the optimal value of (B.35) is 00, and is achievable if and only if $r=r^{*}$. If $r>r^{*}$, then (S1) holds. If $r=r^{*}$, then any feasible solution to (B.35) is optimal. Since $r=r^{*}<m$, for $\hat{d}:=\min\{r,m-r^{*}\}$ and any $w>0$, the point $(w,\hat{d})$ is feasible (because $\min\{r,m-r^{*}\}>0$), and in this case we have $r^{*}-r+\hat{d}=\hat{d}>0$, which means (S2) holds.
    Next we assume $\kappa\neq 3$. Let $\alpha:=\frac{r-r^{*}}{1-\frac{4}{(\kappa-1)^{2}}}=\frac{r-r^{*}}{1-\frac{4\mu^{2}}{(L-\mu)^{2}}}$. We now rewrite $H$ in (B.34) as:
@@
    If $\kappa<3$ and $r\geq r^{*}$, then $\alpha\leq 0$, and we see that the optimal value of (B.35) is 00. Moreover, for all feasible solution $(d,w)$ of (B.35) with $r^{*}-r+d>0$, we have $d>r-r^{*}\geq 0$ and $H(d,w)<0$. Then (S1) holds.
    Finally, we assume $\kappa>3$. If $\alpha>\min\{r,m-r^{*}\}$, then the optimal value of (B.35) is 00, and for all $(d,w)$ that is feasible to (B.35), it holds that $H(d,w)<0$. Then (S1) holds. If $\alpha\leq\min\{r,m-r^{*}\}$, then we can select $d=\min\{r,m-r^{*}\}$ and any $w>0$, which is feasible for (B.35) and $r^{*}-r+d=\min\{r^{*},m-r\}>0$, and we have $H(d,w)\geq 0$. Then (S2) holds.

### v2

@@ 5 Generalization to the RIP case
    where we used the fact that $x_{i}$ and $y_{\tau(i)}$ are 0. Notice that the objective of the above problem is $(g_{i}-v_{\tau(i)})^{2}$. Clearly, the optimal value of (A.18) is 0, and is achieved if and only if
    $g_{i}=v_{\tau(i)}\in[0,1].$ (A.19)
Consequently, by the solution sets given in (A.10), (A.13), (A.16), (A.17) and (A.19), we know the solution set of (A.4) is nonempty. The optimal value is obtained by summing all the optimal values given in the four cases. Moreover, every solution $(\bar{x},\bar{g},\bar{y},\bar{v})$ of (A.4) satisfies $\|\bar{g}\|_{\infty}>1$ if and only if $|\mathcal{J}_{3}^{\tau}|>0$, according to the structure of $\bar{g}$ given in (A.16), (A.17) and (A.19). ∎
Proposition A.3.
Let $r^{*}\in[m]\cup\{0\},r\in[m]$, and $\infty>L\geq\mu>0$. Let $G$ be the objective function of A.3 and let $\kappa:=\frac{L}{\mu}\geq 1$. If $^{*}, rr$ and $\kappa$ satisfy any of the following conditions, then there is no feasible $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ to (A.3) satisfying $\|\bar{g}\|_{\infty}>\lambda$ and $G(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})\geq 0$.
  • (1)
    $r=m$.
  • (2)
    $r\geq r^{*}$ and $\min\{r,m-r^{*}\}>\frac{(\kappa-1)^{2}}{4}\min\{r^{*},m-r\}$.
@@
By the change of variables $(x,g,y,v,\tau,w)\leftarrow(x/\lambda,g/\lambda,y/\lambda,v/\lambda,\tau,w/\lambda)$, we see that A.3 can be reduced to the following problem:
$\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m},w\in\mathbb{R}\end{subarray}}$ $\lambda^{2}\bigg[\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (A.20)
@@
Since dropping the constant $\lambda^{2}$ won’t affect the sign of the function value and our claim only concerns the feasible set of (A.3) and the sign of its objective value, we shall consider the following optimization problem instead:
$\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m},w\in\mathbb{R}\end{subarray}}$ $\sum_{i=1}^{m}(Lx_{i}+g_{i})(\mu y_{\tau(i)}+v_{\tau(i)})+\sum_{i=1}^{m}(\mu x_{i}+g_{i})(Ly_{\tau(i)}+v_{\tau(i)})$ (A.21)
@@
Notice that when $\tau$ and $w$ are fixed, A.21 becomes A.4. Applying Lemma A.2, setting $d_{\tau}=|\mathcal{J}^{\tau}_{2}|$ and recalling the definition of $\mathcal{J}^{\tau}_{3}$ in (A.2), we see that the solution set $\Omega_{w,\tau}$ of (A.4) is nonempty, and for all $(\bar{x},\bar{g},\bar{y},\bar{v})\in\Omega_{w,\tau}$, it holds that $\|\bar{g}\|_{\infty}>1$ if and only if $r^{*}-r+d_{\tau}>0$.
@@
$H(d,w):=\left(-L\mu d+(r^{*}-r+d)\frac{L(L-\mu)^{2}}{4\mu}\right)w^{2}.$ (A.22)
Then in view of (A.2) and (A.5), the optimal value of (A.4) is given by $H(|\mathcal{J}^{\tau}_{2}|,w)$. Moreover, we see that (A.21) is equivalent to the following problem
$\sup_{w\in\mathbb{R},~d\in\mathbb{N}_{0}}H(d,w)\quad s.t.~w>0,~r-r^{*}\leq d\leq\min\{r,m-r^{*}\},$ (A.23)
where the bound for $d$ comes from the requirement that $|\mathcal{J}_{i}^{\tau}|\geq 0$ for $i\in[4]$ in A.2.
@@
    The optimal value of (A.23) is nonpositive, and (A.23) has no feasible solution $(d,w)$ satisfying $H(d,w)\geq 0$ and $r^{*}-r+d>0$.
    In this scenario, we claim that (A.3) has no feasible solution $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ with $\|\bar{g}\|_{\infty}>\lambda$ and $G(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})\geq 0$, where $G$ is the objective of A.3. A short proof is provided below.
    Suppose such a feasible solution $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ of (A.3) exists, then either $(\bar{x},\bar{g},\bar{y},\bar{v})/\lambda$ is optimal for (A.4) with $w=\bar{w}/\lambda$ and $\tau=\bar{\tau}$, or $(\bar{x},\bar{g},\bar{y},\bar{v})/\lambda$ is not optimal. In the latter case, the optimal value of (A.23) must be positive. In the former case, we see from Lemma A.2 that $|\mathcal{J}_{3}^{\bar{\tau}}|>0$, and hence (A.23) has a feasible solution $(\tilde{d},\tilde{w})=(|\mathcal{J}^{\bar{\tau}}_{2}|,\bar{w}/\lambda)$ with $H(\tilde{d},\tilde{w})\geq 0$ and $r^{*}-r+\tilde{d}=r^{*}-r+|\mathcal{J}^{\bar{\tau}}_{2}|=|\mathcal{J}^{\bar{\tau}}_{3}|>0$ (see (A.2)). Both cases yield a contradiction.
@@
[…]
We note that the classification in (S1) and (S2) is not complete, since we cannot say anything if the optimal value of (A.23) is positive and there is no feasible solution $(d,w)$ of (A.23) satisfying $H(d,w)\geq 0$ and $r^{*}-r+d>0$. Nevertheless, the two scenarios in (S1) and (S2) are enough for our proof. Consider the following cases on $r$, $r^{*}$ and $\kappa:=L/\mu$.
@@
    $r=m$ or $r^{*}=0$. If $r=m$, then every feasible point $(d,w)$ to (A.23) must satisfy $d=r-r^{*}$ and $H(d,w)=-L\mu dw^{2}\leq 0$. Hence, (S1) holds. If $r^{*}=0$, we see that every feasible point $(d,w)$ to (A.23) must satisfy $d=r$. Then, in view of (A.22), we can rewrite (A.23) as:
@@
    This means that every feasible solution of (A.23) has a negative objective function value. Then (S1) holds.
@@
    Then the optimal value of (A.23) is 0, and for all feasible solution $(d,w)$ of (A.23) with $r^{*}-r+d>0$ it holds that $d>r-r^{*}\geq 0$, and hence $H(d,w)<0$. Then (S1) holds.
@@
    $m>r\geq r^{*}>0$ and $L>\mu$ (i.e., $\kappa>1$). If $\kappa=\frac{L}{\mu}=3$, then we have $L\mu+\frac{L(L-\mu)^{2}}{4\mu}=L\mu(\frac{(\kappa-1)^{2}}{4}-1)=0$. Therefore, (A.23) is equivalent to that:
@@
[…]
    Next we assume $\kappa\neq 3$. Let $\alpha:=\frac{r-r^{*}}{1-\frac{4}{(\kappa-1)^{2}}}=\frac{r-r^{*}}{1-\frac{4\mu^{2}}{(L-\mu)^{2}}}$. We now rewrite $H$ in (A.22) as:
@@
    If $\kappa<3$ and $r\geq r^{*}$, then $\alpha\leq 0$, and we see that the optimal value of (A.23) is 0. Moreover, for all feasible solution $(d,w)$ of (A.23) with $r^{*}-r+d>0$, we have $d>r-r^{*}\geq 0$ and $H(d,w)<0$. Then (S1) holds.
    Finally, we assume $\kappa>3$. If $\alpha>\min\{r,m-r^{*}\}$, then the optimal value of (A.23) is 0, and for all $(d,w)$ that is feasible to (A.23), it holds that $H(d,w)<0$. Then (S1) holds. If $\alpha\leq\min\{r,m-r^{*}\}$, then we can select $d=\min\{r,m-r^{*}\}$ and any $w>0$, which is feasible for (A.23) and $r^{*}-r+d=\min\{r^{*},m-r\}>0$, and we have $H(d,w)\geq 0$. Then (S2) holds.